# Bayesian Time-Varying Meta-Analysis via Hierarchical Mean-Variance Random-effects Models

### v1

@@ Introduction
Meta-analysis is used to integrate results from multiple experiments to obtain generalized insights in fields such as epidemiology (Kleinbaum et al., 1991), economics (Stanley and Doucouliagos, 2012), and engineering (Sohn, 1999). In meta-analyses, heterogeneous effects among experiments are modeled using meta-regression. To address between-study heterogeneity from subgroups with different true effects, hierarchical multilevel models (e.g. Pastor and Lazowski, 2018), particularly hierarchical Bayesian models incorporating random effects (e.g. Higgins et al., 2009; Browne and Jones, 2017), are commonly used. These models typically assume homoscedasticity across datasets (e.g. Higgins et al., 2009).
@@
Real-world meta-analysis datasets often involve multiple experiments with time-varying properties, leading to high heteroscedasticity and temporal heterogeneity. For example, promotional marketing experiments on digital platforms, such as those of retailers and media providers, exhibit such heterogeneity by conducting experiments with diverse participants and media at different times. These situations result in different experimental means and variances, even among similar promotions. Standard random-effects meta-analyses assume homoscedasticity and neglect temporal heterogeneity, potentially inadequately addressing these issues. Therefore, it is necessary to model the high heteroscedasticity and time trends appropriately.
@@
Several approaches have been proposed to address heteroscedasticity. Thompson and Becker (2020) introduced weakly informative priors for subgroup-specific standard deviations. Bowater and Escarela (2013) suggested making variance proportional to experiment-specific variables, such as participant numbers. Williams et al. (2021) developed a Bayesian fixed-effect model using experiment-related variables for each experiment’s variance. Nevertheless, these methods can lead to unreliable estimates due to large standard errors in small-scale experiments with limited sample sizes. In meta-analysis literature addressing time effects, some methods compare time-varying functions across experiments using dose-response and time-response models (Mawdsley et al., 2016; Pedder et al., 2019). Others treat experiment timing as fixed effects (Miller and Hosanagar, 2020; Salanti et al., 2009; van Enst et al., 2015). However, the former does not account for treatment effects changing over time, limiting its applicability in scenarios with time-varying effects among experiments. The latter uses discrete dummy variables for time effects, leading to unstable estimates when time effects vary continuously. Consequently, these approaches can be unsuitable for high heteroscedasticity and temporal heterogeneity.
@@
To solve the aforementioned problems, we propose a new Bayesian estimation method called Shrinkage estimation with Time-trend and Random-Effects for meta-Analysis of Means and variance (STREAM). STREAM comprises two components: shrinkage estimation of means and variances through a hierarchical Bayesian approach that applies “borrowing strength” and estimation of time effects using the Gaussian process to flexibly model continuous-time trends in datasets.
@@
We formulate the problem of meta-analysis for estimating experimental effects using a collected experimental dataset. This dataset includes $m$ experiments, and let $\theta_{i}$ for $i=1,2,\ldots,m$ denote the experiment-level parameters of interest. The observed data for each experiment is an estimate $y_{i}$ of $\theta_{i}$ and its associated variance $S_{i}^{2}$, based on a sample with size $n_{i}$.
@@
For the observed data, we assume that
$\begin{split}&y_{i}\mid\theta_{i},\sigma_{i}^{2}\sim N(\theta_{i},\sigma_{i}^{2}),\\ &S_{i}^{2}\mid\sigma_{i}^{2}\sim\Gamma\left(\frac{n_{i}-1}{2},\frac{n_{i}-1}{2\sigma_{i}^{2}}\right),\\ \end{split}$ (1)
where $\Gamma(\alpha,\beta)$ represents the gamma distribution expressed as $x^{(\alpha-1)}\exp(-\beta x),x>0$. Here $\sigma_{i}^{2}$ is the true sampling variance, representing the variability of $y_{i}$ as an estimate of $\theta_{i}$. The parameter $\sigma_{i}^{2}$ is unknown, and the observed $S_{i}^{2}$ is an estimate of $\sigma_{i}^{2}$. The distributional assumption for $S_{i}^{2}$ in (1) is equivalent to assuming that $S_{i}^{2}/\sqrt{\sigma_{i}^{2}/n_{i}}$ follows a $\chi^{2}$-distribution with $(n_{i}-1)$ degrees of freedom. Hence, the distributional assumptions for $y_{i}$ and $S_{i}^{2}$ in (1) are reasonable as long as $n_{i}$ is not too small. Further, it would also be reasonable to assume that $y_{i}$ and $S_{i}^{2}$ are conditionally independent given $\theta_{i}$ and $\sigma_{i}^{2}$.
@@
$\begin{split}&\theta_{i}=\alpha_{\theta}+\bm{\theta}_{a}^{\top}\bm{z}_{ai}+\bm{\theta}_{b}^{\top}\bm{z}_{bi}+\bm{\theta}_{c}^{\top}\bm{z}_{ci}+\bm{\beta}_{\theta}^{\top}\bm{x}_{i},\\ &\theta_{aj}\sim N(0,\tau_{a}^{2}),\ \ j=1,\ldots,J,\ \ \ \theta_{bk}\sim N(0,\tau_{b}^{2}),\ \ k=1,\ldots,K,\\ &\bm{\theta}_{c}\sim N(0_{L},C(\sigma_{p}^{2},l_{p})),\\ \end{split}$ (2)
where $\alpha_{\theta}$ represents the baseline value common to each experiment, $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ denote the random effects for the groups to which each experiment belongs, expressed as $J$-dimensional and $K$-dimensional vectors, respectively. Moreover, $\bm{\theta}_{c}$ represents the time effect in experiment $i$. $\bm{\beta}_{\theta}$ represents the regression coefficients for the experiment characteristics $\bm{x}_{i}$. All elements of $\bm{\theta}_{a}=(\theta_{a1},\ldots,\theta_{aJ})$, $\bm{\theta}_{b}=(\theta_{b1},\ldots,\theta_{bK})$ and $\bm{\theta}_{c}=(\theta_{c1},\ldots,\theta_{cL})$ are assumed to be independent, and $\alpha$, $\bm{\beta}_{\theta}$, $\tau_{a}^{2}$ and $\tau_{b}^{2}$ are unknown parameters. Here $C(\sigma_{p}^{2},l_{p})$ is the $L\times L$ variance-covariance matrix induced by the Gaussian process assumption on $\bm{\theta}_{c}$ as a function $t_{i}$, and its $(l,l^{\prime})$-element is $K(t_{l},t_{l^{\prime}};\sigma_{p}^{2},l_{p})$ for some kernel function with parameters $\sigma_{p}^{2}$ and $l_{p}$. The kernel function, which determines the covariance structures of the Gaussian process, has various options. Since the data we will consider in our application contains time information as date, we use the following periodic covariance kernel (Roberts et al., 2013; Williams and Rasmussen, 2006) to take account of annual periodicity:
$\begin{split}&K(t,t^{\prime};\sigma_{p}^{2},l_{p})=\sigma_{p}^{2}\exp{\left(-\frac{2\sin^{2}{(\pi|t-t^{\prime}|/p_{e})}}{l_{p}^{2}}\right)},\\ \end{split}$
@@
$\begin{split}&\log\sigma_{i}^{2}\sim N(\alpha_{\sigma}+\bm{\delta}_{a}^{\top}\bm{z}_{ai}+\bm{\delta}_{b}^{\top}\bm{z}_{bi}+\bm{\beta}_{\sigma}^{\top}\bm{x_{i}}-\log n_{i},\tau_{\sigma}^{2}),\\ &\delta_{aj}\sim N(0,\tau_{c}^{2}),\ \ j=1,\ldots,J,\ \ \ \delta_{bk}\sim N(0,\tau_{d}^{2}),\ \ k=1,\ldots,K,\\ \end{split}$ (3)
where $\alpha_{\sigma}$ represents the baseline value common to each experiment, $\bm{\delta}_{a}$ and $\bm{\delta}_{b}$ denote the random effects for the groups to which each experiment belongs, expressed as $J$-dimensional and $K$-dimensional vectors, respectively, and $\bm{\beta}_{\sigma}$ represents the regression coefficients for the experiment characteristics $\bm{x}_{i}$. Further, $\alpha_{\sigma}$, $\bm{\beta}_{\sigma}$, $\tau_{\sigma}^{2}$, $\tau_{c}^{2}$ and $\tau_{d}^{2}$ are unknown parameters. Note that $-\log n_{i}$ in the mean term of (3) corresponds to an offset term to ensure that $E[\sigma_{i}^{2}]=O(n_{i}^{-1})$. The proposed model consists of (1), (2) and (3), and its advantages can be summarized as follows:
@@
We calculate the predictive distribution of the estimates $\tilde{\bm{y}}=(\tilde{y}_{1},\ldots,\tilde{y}_{m_{\text{te}}})$ for $m_{\text{te}}$ new future experiments. Specifically, we use $m_{\text{tr}}$ experiments as observational data and the posterior samples of each parameter generated by MCMC. In constructing the predictive distribution of $\tilde{\bm{y}}$, we use the posterior samples generated by MCMC for various parameters in equations (1), (2), and (3). We provide additional explanation since constructing the predictive distribution for $\bm{\theta}_{c}$ differs from other parameters. Our goal is to construct the predictive distribution of the parameters $\tilde{\bm{\theta}}_{c}=(\tilde{\theta}_{c1},\ldots,\tilde{\theta}_{cm_{\text{te}}})$ associated with the future experiments $\tilde{\bm{z}}_{c}=(\tilde{\bm{z}}_{c1},\ldots,\tilde{\bm{z}}_{cm_{\text{te}}})$, using the posterior samples. To achieve this, we aim to obtain the predictive distribution $p(\tilde{\bm{\theta}}_{c}|D_{\text{tr}},\tilde{\bm{z}}_{c})$. Here, $D_{\text{tr}}=(\bm{z}_{c},\bm{\theta}_{c})$, where $\bm{z}_{c}=(z_{c1},\ldots,z_{cm_{\text{tr}}})$ and $\bm{\theta}_{c}=(\theta_{c1},\ldots,\theta_{cm_{\text{tr}}})$. Since the joint distribution $p(\tilde{\bm{\theta}}_{c},\bm{\theta}_{c}|\bm{z}_{c},\tilde{\bm{z}}_{c})$ is given by the Gaussian process assumption in equation (2), it is formulated as the following multivariate Gaussian distribution.
$\begin{split}p(\bm{\theta}_{c},\tilde{\bm{\theta}}_{c}|\bm{z}_{c},\tilde{\bm{z}}_{c})\sim N(0_{m_{\text{tr}}+m_{\text{te}}},C^{\prime}(\sigma^{2}_{p},l_{p}))\end{split}$ (5)
where $C^{\prime}(\sigma_{p}^{2},l_{p})$ is the $L\times L$ the following variance-covariance matrix:
$\begin{split}C^{\prime}(\sigma^{2}_{p},l_{p})=\begin{pmatrix}\bm{\Phi}_{m_{\text{tr}}m_{\text{tr}}}&\bm{\Phi}_{m_{\text{tr}}m_{\text{te}}}\\ \bm{\Phi}^{\top}_{m_{\text{tr}}m_{\text{te}}}&\bm{\Phi}_{m_{\text{te}}m_{\text{te}}}\end{pmatrix}\end{split}$ (6)
@@
$\begin{split}&\bm{\Phi}_{m_{\text{tr}}m_{\text{tr}}}=(K(t_{i},t_{j};\sigma^{2}_{p},l_{p})),\ \ 1\leq i,j\leq m_{\text{tr}},\\ &\bm{\Phi}_{m_{\text{tr}}m_{\text{te}}}=(K(t_{i},\tilde{t}_{j};\sigma^{2}_{p},l_{p})),\ \ 1\leq i\leq m_{\text{tr}},\ \ 1\leq j\leq m_{\text{te}}\\ &\bm{\Phi}_{m_{\text{te}}m_{\text{te}}}=(K(\tilde{t}_{i},\tilde{t}_{j};\sigma^{2}_{p},l_{p})),\ \ 1\leq i,j\leq m_{\text{te}},\\ \end{split}$ (7)
@@
$\begin{split}p(\tilde{\bm{\theta}}_{c}|D_{tr},\tilde{\bm{z}}_{c})=N(\bm{\Phi}^{\top}_{m_{\text{tr}}m_{\text{te}}}\bm{\Phi}^{-1}_{m_{\text{tr}}m_{\text{tr}}}\bm{\theta}_{c},\bm{\Phi}_{m_{\text{te}}m_{\text{te}}}-\bm{\Phi}^{\top}_{m_{\text{tr}}m_{\text{te}}}\bm{\Phi}^{-1}_{m_{\text{tr}}m_{\text{tr}}}\bm{\Phi}_{m_{\text{tr}}m_{\text{te}}}).\end{split}$ (8)
Then, using the posterior samples of $\alpha_{\theta},\bm{\theta}_{a},\bm{\theta}_{b}$ and $\bm{\beta}_{\theta}$, the posterior samples of the future effect $\tilde{\theta}_{j}$ can be generated as
@@
$\begin{split}&y_{ij}\sim N(\theta_{i},\sigma_{i}^{2}),\ \ \ i=1,\ldots,m,\ j=1,\ldots,n_{i},\\ \end{split}$ (9)
@@
$\begin{split}&\theta_{i}=\bm{\theta}_{a}^{\top}\bm{z}_{ai}+\bm{\theta}_{b}^{\top}\bm{z}_{bi}+a_{1}\sin\left(2\pi\frac{t_{i}}{12}\right)+b_{1}\cos\left(2\pi\frac{t_{i}}{12}\right)+c_{1}\frac{t_{i}}{12}+0.5x_{i},\\ &\theta_{aj},\theta_{bk}\sim N(\mu_{\theta},\sigma_{\theta}^{2}),\ \ \ \mu_{\theta}\sim N(3,2^{2}),\ \ \ \sigma_{\theta}^{2}\sim N_{[0,\infty)}(0,d_{1}^{2}),\\ \end{split}$ (11)
@@
For each simulation scenario, we generate 150 simulation datasets and estimate the experimental effect $\theta_{i}$ using seven baseline methods and the proposed STREAM: (a) Fixed Effects (FE) model, (b) FE with monthly dummy (FE-M) model, (c) FE with monthly dummy and variance modeling (FE-MV) model, (d) Random Effects (RE) model, (e) RE with monthly dummy (RE-M) model, (f) RE with monthly dummy and variance modeling (RE-MV) model, (g) RE-Gaussian-Process (RE-GP) model, and (h) STREAM. The FE model, corresponding to the fixed effects model in general meta-analysis (Borenstein et al., 2010), models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ as fixed effects in equation (2) and does not account for the time effect $\bm{\theta}_{c}$, using an uninformative prior distribution $N(0,1000)$ for $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$. The FE-M model extends the FE model by including the time effect $\bm{\theta}_{c}$ as a fixed effect, with an uninformative prior $N(0,1000)$ for $\bm{\theta}_{c}$. The FE-MV model further incorporates fixed effects for $\bm{\delta}_{a}$ and $\bm{\delta}_{b}$ in equation (3), also using an uninformative prior $N(0,1000)$. This model corresponds to existing methods by Bowater and Escarela (2013) and Williams et al. (2021). The RE model, corresponding to hierarchical Bayesian random effects modeling in general meta-analysis (Babapulle et al., 2004; Borenstein et al., 2010), models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ as random effects in equation (2) without considering the time effect $\bm{\theta}_{c}$. The RE-M model extends the RE model by modeling the time effect $\bm{\theta}_{c}$ as a fixed effect with an uninformative prior $N(0,1000)$. The RE-MV model further incorporates random effects for $\bm{\delta}_{a}$ and $\bm{\delta}_{b}$ in equation (3). The RE-GP model models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ as random effects in equation (2) and represents the time effect $\bm{\theta}_{c}$ using a Gaussian process, corresponding to the method in LeBlanc and Banks (2022). For the baseline methods FE, RE, FE-M, RE-M, and RE-GP that do not model variance, the estimated variance uses the average variance obtained from the training dataset as a fixed variance.
@@
scaled $\text{scaled MSE}=\frac{1}{n}\sum_{i=1}^{n}\frac{(\theta_{i}-\hat{y}_{i})^{2}}{\theta_{i}^{2}},$ (15)
@@
We use the interval score (IS) (Gneiting and Raftery, 2007) to evaluate the interval estimation performance of the models. Specifically, the IS for the $100(1-\alpha)$prediction interval, with lower and upper endpoints that are the predictive quantiles at level $\alpha/2$ and $(1-\alpha)/2$, respectively, are defined below.
@@
where $l_{i}$ and $u_{i}$ are the lower and upper bounds of the $100(1-\alpha)$highest posterior density intervals, respectively, and $\mathbbm{1}(\cdot)$ is the indicator function. The IS is a measure designed to quantify the quality of uncertainty in the model predictions. It penalizes the model if the prediction interval it outputs is vast and penalizes the model if the observed values are outside the prediction interval it outputs. In this paper, $\alpha$ is 0.05.
@@
Table 5: Median and maximum values of Gelman-Rubin (GR) statistic $\hat{R}$. The boldface values indicate $\hat{R}<1.1$, satisfying the convergence criterion for MCMC.

### v2

@@ Introduction
Meta-analysis is used to integrate results from multiple experiments to obtain generalized insights in fields such as epidemiology (Kleinbaum et al., 1991), economics (Stanley and Doucouliagos, 2012), and engineering (Sohn, 1999). In meta-analyses, heterogeneous effects among experiments are modeled using meta-regression. To address between-study heterogeneity from subgroups with different true effects, hierarchical multilevel models (e.g. Pastor and Lazowski, 2018), particularly hierarchical Bayesian models incorporating random effects (e.g. Higgins et al., 2009; Browne and Jones, 2017), are commonly used. While these models typically allow individual study effects to vary, they generally consolidate what could be multiple, distinct between-study variances into a single summary parameter (e.g., $\tau^{2}$). This approach inherently assumes that this single, representative level of dispersion is constant throughout the meta-analysis (e.g. Higgins et al., 2009; Williams et al., 2021).
@@
A central challenge in real-world meta-analysis of promotional marketing experiments is predicting future experimental effects from accumulated past results. This challenge is complicated by datasets often involving multiple experiments with time-varying properties, leading to high heteroscedasticity and temporal heterogeneity. For example, in marketing, promotional experiments on digital platforms by retailers and media providers often exhibit such heterogeneity because they are conducted with diverse participants and media at different times. These conditions result in varying experimental means and variances, even among similar promotions. Consequently, predicting future experimental effects from these complex accumulated datasets is crucial for designing effective future promotional strategies. Since promotional effects can exhibit seasonality or other periodic patterns, capturing these periodic structures within a predictive framework can be particularly effective for accurate forecasting. Standard random-effects meta-analyses, however, typically assume homoscedasticity and neglect temporal heterogeneity, potentially inadequately addressing these complexities and thus limiting their predictive performance. Therefore, it is necessary to appropriately model high heteroscedasticity and complex time trends, including periodic components, to enhance understanding and predictive performance in such contexts.
@@
Several approaches have been proposed to address high heteroscedasticity including between-study variances and complex time trends. Thompson and Becker (2020) introduced weakly informative priors for subgroup-specific standard deviations. Bowater and Escarela (2013) suggested making variance proportional to experiment-specific variables, such as participant numbers. Williams et al. (2021) developed a Bayesian fixed-effect model using experiment-related variables for each experiment’s variance. Nevertheless, these methods can lead to unreliable estimates due to large standard errors in small-scale experiments with limited sample sizes. In meta-analysis literature addressing time effects, some methods compare time-varying functions across experiments using dose-response and time-response models (Mawdsley et al., 2016; Pedder et al., 2019). Others treat experiment timing as fixed effects (Miller and Hosanagar, 2020; Salanti et al., 2009; van Enst et al., 2015). However, the former does not account for treatment effects changing over time, limiting its applicability in scenarios with time-varying effects among experiments. The latter uses discrete dummy variables for time effects, leading to unstable estimates when time effects vary continuously. Consequently, these approaches can be unsuitable for high heteroscedasticity and temporal heterogeneity.
@@
To solve the aforementioned problems, we propose a new Bayesian estimation method called Shrinkage estimation with Time-trend and Random-Effects for meta-Analysis of Means and variance (STREAM). STREAM comprises two components: First, shrinkage estimation of means and variances through a hierarchical Bayesian approach that applies “borrowing strength”. In this component, while estimates of variances are produced, the principal motivation for robustly modeling these variances is to improve the performance and reliability of the mean effect estimation. Second, estimation of time effects using the Gaussian process to flexibly model continuous-time trends in datasets.
@@
We formulate the problem of meta-analysis for estimating experimental effects from a collection of $m$ individual experiments. For each experiment $i$ ($i=1,\dots,m$), let $\theta_{i}$ denote the true underlying effect or parameter of interest. The observed data typically consist of $y_{i}$, which is an estimate of $\theta_{i}$, and an associated measure of its precision. Specifically, $\sigma_{i}^{2}$ is defined as the true unknown sampling variance of the estimate $y_{i}$. The data available from each experiment $i$, based on a sample of size $n_{i}$, includes $y_{i}$ and $S_{i}^{2}$, where $S_{i}^{2}$ is an observed estimate of this true sampling variance $\sigma_{i}^{2}$.
To build intuition, consider a common scenario where $y_{i}$ is the sample mean of $n_{i}$ independent observations drawn from a population with true mean $\theta_{i}$ and true variance $\psi_{i}^{2}$ (representing the variance of an individual observation in study $i$). In this case, the true sampling variance of $y_{i}$ (which we denote as $\sigma_{i}^{2}$ throughout this paper) is given by $\sigma_{i}^{2}=\psi_{i}^{2}/n_{i}$, and this $\sigma_{i}^{2}$ is of order $n_{i}^{-1}$. The observed $S_{i}^{2}$ would then be an estimate of this $\sigma_{i}^{2}$, typically calculated as $s_{i}^{2}/n_{i}$, where $s_{i}^{2}$ is the sample variance of the $n_{i}$ observations (an estimate of $\psi_{i}^{2}$).
In conventional meta-analysis, the observed estimates $S_{i}^{2}$ are often treated as if they were the true sampling variances $\sigma_{i}^{2}$ and are used as fixed, known weights (typically the inverse of these variances) in pooling procedures. Some simpler approaches might even disregard individual $S_{i}^{2}$ values if they assume a common variance across studies, though this is less common when study-specific variance estimates are available. In contrast, our proposed model explicitly acknowledges $S_{i}^{2}$ as an estimate of an unknown $\sigma_{i}^{2}$ and models the relationship and uncertainty, as detailed in Equation (1).
@@
For the observed data $(y_{i},S_{i}^{2})$ from each experiment $i$, we assume the following distributions:
$\begin{split}y_{i}\mid\theta_{i},\sigma_{i}^{2}&\sim N(\theta_{i},\sigma_{i}^{2}),\\ S_{i}^{2}\mid\sigma_{i}^{2}&\sim\Gamma\left(\frac{n_{i}-1}{2},\frac{n_{i}-1}{2\sigma_{i}^{2}}\right).\end{split}$ (1)
Here, $\Gamma(\alpha,\beta)$ denotes the gamma distribution with a probability density function proportional to $x^{\alpha-1}\exp(-\beta x)$ for $x>0$. The distributional assumption for $S_{i}^{2}$ in (1) is standard and implies that $(n_{i}-1)S_{i}^{2}/\sigma_{i}^{2}$ follows a $\chi^{2}$-distribution with $n_{i}-1$ degrees of freedom. These assumptions for $y_{i}$ and $S_{i}^{2}$ are generally considered reasonable, particularly when the sample size $n_{i}$ for each study is not excessively small. Furthermore, we assume that $y_{i}$ and $S_{i}^{2}$ are conditionally independent given their respective underlying parameters $\theta_{i}$ and $\sigma_{i}^{2}$.
@@
[…]
where $\alpha_{\theta}$ represents the baseline value common to each experiment, $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ denote the random effects for the groups to which each experiment belongs, expressed as $J$-dimensional and $K$-dimensional vectors, respectively. Moreover, $\bm{\theta}_{c}$ represents the time effect. $\bm{\beta}_{\theta}$ represents the regression coefficients for the experiment characteristics $\bm{x}_{i}$. All elements of $\bm{\theta}_{a}=(\theta_{a1},\ldots,\theta_{aJ})$ and $\bm{\theta}_{b}=(\theta_{b1},\ldots,\theta_{bK})$ are assumed to be independent, and $\alpha$, $\bm{\beta}_{\theta}$, $\tau_{a}^{2}$ and $\tau_{b}^{2}$ are unknown parameters. Here $C(\sigma_{p}^{2},l_{p})$ is the $L\times L$ variance-covariance matrix induced by the Gaussian process assumption on $\bm{\theta}_{c}$ as a function of $t_{i}$, and its $(l,l^{\prime})$-element is $K(t_{l},t_{l^{\prime}};\sigma_{p}^{2},l_{p})$ for some kernel function with parameters $\sigma_{p}^{2}$ and $l_{p}$. The kernel function, which determines the covariance structures of the Gaussian process, has various options. Since the data we will consider in our application contains time information as date, we use the following periodic covariance kernel (Roberts et al., 2013; Williams and Rasmussen, 2006) to take account of annual periodicity:
[…]
@@
[…]
where $\alpha_{\sigma}$ represents the baseline value common to each experiment, $\bm{\delta}_{a}$ and $\bm{\delta}_{b}$ denote the random effects for the groups to which each experiment belongs, expressed as $J$-dimensional and $K$-dimensional vectors, respectively, and $\bm{\beta}_{\sigma}$ represents the regression coefficients for the experiment characteristics $\bm{x}_{i}$. Further, $\alpha_{\sigma}$, $\bm{\beta}_{\sigma}$, $\tau_{\sigma}^{2}$, $\tau_{c}^{2}$ and $\tau_{d}^{2}$ are unknown parameters. Note that $\log n_{i}$ in the mean term of (3) corresponds to an offset term to ensure that $E[\sigma_{i}^{2}]=O(n_{i}^{-1})$. The proposed model consists of (1), (2) and (3), and its advantages can be summarized as follows:
@@
We calculate the predictive distribution of the estimates $\tilde{\bm{y}}=(\tilde{y}_{1},\ldots,\tilde{y}_{m_{\text{te}}})$ for $m_{\text{te}}$ new future experiments with specific time point from the set $\{t^{\prime}_{1},\ldots,t^{\prime}_{L^{\prime}}\}$. Specifically, we use $m_{\text{tr}}$ experiments as observational data and the posterior samples of each parameter generated by MCMC. In constructing the predictive distribution of $\tilde{\bm{y}}$, we use the posterior samples generated by MCMC for various parameters in equations (1), (2), and (3). We provide additional explanation since constructing the predictive distribution for $\bm{\theta}_{c}$ differs from other parameters. Our goal is to construct the predictive distribution of the parameters $\tilde{\bm{\theta}}_{c}=(\tilde{\theta}_{c1},\ldots,\tilde{\theta}_{cL^{\prime}})$ associated with the future experiments $\tilde{\bm{z}}_{c}=(\tilde{\bm{z}}_{c1},\ldots,\tilde{\bm{z}}_{cL^{\prime}})$, using the posterior samples. To achieve this, we aim to obtain the predictive distribution $p(\tilde{\bm{\theta}}_{c}|D_{\text{tr}},\tilde{\bm{z}}_{c})$. Here, $D_{\text{tr}}=(\bm{z}_{c},\bm{\theta}_{c})$, where $\bm{z}_{c}=(\bm{z}_{c1},\ldots,\bm{z}_{cL})$ and $\bm{\theta}_{c}=(\theta_{c1},\ldots,\theta_{cL})$. Since the joint distribution $p(\tilde{\bm{\theta}}_{c},\bm{\theta}_{c}|\bm{z}_{c},\tilde{\bm{z}}_{c})$ is given by the Gaussian process assumption in equation (2), it is formulated as the following multivariate Gaussian distribution.
$\begin{split}p(\bm{\theta}_{c},\tilde{\bm{\theta}}_{c}|\bm{z}_{c},\tilde{\bm{z}}_{c})\sim N(0_{L+L^{\prime}},C^{\prime}(\sigma^{2}_{p},l_{p}))\end{split}$ (5)
where $C^{\prime}(\sigma_{p}^{2},l_{p})$ is the $(L+L^{\prime})\times(L+L^{\prime})$ the following variance-covariance matrix:
$\begin{split}C^{\prime}(\sigma^{2}_{p},l_{p})=\begin{pmatrix}\bm{\Phi}_{LL}&\bm{\Phi}_{LL^{\prime}}\\ \bm{\Phi}^{\top}_{LL^{\prime}}&\bm{\Phi}_{L^{\prime}L^{\prime}}\end{pmatrix}\end{split}$ (6)
@@
[…]
@@
$\begin{split}p(\tilde{\bm{\theta}}_{c}|D_{tr},\tilde{\bm{z}}_{c})=N(\bm{\Phi}^{\top}_{LL^{\prime}}\bm{\Phi}^{-1}_{LL}\bm{\theta}_{c},\bm{\Phi}_{L^{\prime}L^{\prime}}-\bm{\Phi}^{\top}_{LL^{\prime}}\bm{\Phi}^{-1}_{LL}\bm{\Phi}_{LL^{\prime}}).\end{split}$ (8)
Then, using the posterior samples of $\alpha_{\theta},\bm{\theta}_{a},\bm{\theta}_{b},\tilde{\bm{\theta}}_{c}$ and $\bm{\beta}_{\theta}$, the posterior samples of the future effect $\tilde{\theta}_{j}$ can be generated as
@@
[…]
@@
[…]
@@
For each simulation scenario, we generate 150 simulation datasets and estimate the experimental effect $\theta_{i}$ using seven baseline methods and the proposed STREAM: (a) Fixed Effects (FE) model, (b) FE with monthly dummy (FE-M) model, (c) FE with monthly dummy and variance modeling (FE-MV) model, (d) Random Effects (RE) model, (e) RE with monthly dummy (RE-M) model, (f) RE with monthly dummy and variance modeling (RE-MV) model, (g) RE-Gaussian-Process (RE-GP) model, and (h) STREAM. The FE model, corresponding to the fixed effects model in general meta-analysis (Borenstein et al., 2010), models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ as fixed effects in equation (2) and does not account for the time effect $\bm{\theta}_{c}$, using an uninformative prior distribution $N(0,1000)$ for $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$. The FE-M model extends the FE model by including the time effect $\bm{\theta}_{c}$ as a fixed effect, with an uninformative prior $N(0,1000)$ for $\bm{\theta}_{c}$. The FE-MV model further incorporates fixed effects for $\bm{\delta}_{a}$ and $\bm{\delta}_{b}$ in equation (3), also using an uninformative prior $N(0,1000)$. This model corresponds to existing methods by Bowater and Escarela (2013) and Williams et al. (2021). The RE model, corresponding to hierarchical Bayesian random effects modeling in general meta-analysis (Babapulle et al., 2004; Borenstein et al., 2010), models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ as random effects in equation (2) without considering the time effect $\bm{\theta}_{c}$. The RE-M model extends the RE model by modeling the time effect $\bm{\theta}_{c}$ as a fixed effect with an uninformative prior $N(0,1000)$. The RE-MV model further incorporates random effects for $\bm{\delta}_{a}$ and $\bm{\delta}_{b}$ in equation (3). The RE-GP model models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ as random effects in equation (2) and represents the time effect $\bm{\theta}_{c}$ using a Gaussian process, corresponding to the method in LeBlanc and Banks (2022).
These baseline methods adopt distinct approaches for incorporating and modeling $S_{i}^{2}$. Broadly, they are classified based on whether they explicitly model variance heterogeneity. For the first group of baseline methods–FE, RE, FE-M, RE-M, and RE-GP–which do not explicitly model variance heterogeneity, the observed variance estimate for each study $i$, $S_{i}^{2}$, is treated as a known and fixed value representing the sampling variance $\sigma_{i}^{2}$ for that study. Consequently, each study’s observed effect $y_{i}$ is assumed to follow a normal distribution with mean $\theta_{i}$ and this fixed variance $S_{i}^{2}$. Subsequently, when predicting the effect of new experiments for which individual variances are unknown, these baseline methods use a single, fixed variance for all such predictions; this fixed variance is calculated as the average of the observed $S_{i}^{2}$ values from the training dataset.
In contrast, the FE-MV and RE-MV explicitly model the observed sample variance $S_{i}^{2}$ conditional on the true sampling variance $\sigma_{i}^{2}$ using a Gamma distribution, as specified in Equation 1. This modeling strategy for $S_{i}^{2}\mid\sigma_{i}^{2}$ is equivalent to the standard statistical assumption that $(n_{i}-1)S_{i}^{2}/\sigma_{i}^{2}$, follows a Chi-squared distribution. Such an approach to modeling sample variances via a Gamma distribution has precedent in the meta-analysis literature (e.g. Chowdhry et al., 2016; Dakin et al., 2011).
@@
$\text{scaled MSE}=\frac{1}{n}\sum_{i=1}^{n}\frac{(\theta_{i}-\hat{y}_{i})^{2}}{\theta_{i}^{2}},$ (15)
@@
We use the interval score (IS) (Gneiting and Raftery, 2007) to evaluate the interval estimation performance of the models. Specifically, the IS for the $100(1-\alpha)$% prediction interval, with lower and upper endpoints that are the predictive quantiles at level $\alpha/2$ and $(1-\alpha)/2$, respectively, are defined below.
@@
where $l_{i}$ and $u_{i}$ are the lower and upper bounds of the $100(1-\alpha)$% highest posterior density intervals, respectively, and $\mathbbm{1}(\cdot)$ is the indicator function. The IS is a measure designed to quantify the quality of uncertainty in the model predictions. It penalizes the model if the prediction interval it outputs is vast and penalizes the model if the observed values are outside the prediction interval it outputs. In this paper, $\alpha$ is 0.05.
@@
Table 5: Median and maximum values of Gelman-Rubin (GR) statistic $\hat{R}$. The boldface values indicate $\hat{R}<1.1$, satisfying the convergence criterion for $MC$.